OpenStudy (anonymous):

x(1+x)=6 why is it no solution?

OpenStudy (anonymous):

x=2 works

OpenStudy (anonymous):

directions: find all real solutions if any idk why the answer is "no solutions" in the back of the textbook

OpenStudy (noelgreco):

The answers are written by bored graduate students who receive little or no pay. There is no penalty if they are wrong once in a while.

OpenStudy (noelgreco):

x^2 + x - 6 = 0 (x + 3)(x - 2)=0

OpenStudy (anonymous):

oh. okay, thanks